A uniform rod of mass 200g and length,L is supported at both ends.Masses of 300g and 500g are suspended at L/3 and 3L/4 from one end. Determine the force exerted on each support.
1
Expert's answer
2020-10-02T07:17:06-0400

From the principle of moments,

sum of upward forces = sum of downward forces

moment "=" force("F") "\times" distance("L")

moment "=" "F \times L"

Taking moment about one end A we have,

moment "=" "\frac{500}{100} \times \frac{3L}{4} + \frac{200}{100} \times \frac{1L}{2} + \frac{300}{100} \times \frac{1L}{3}"

"=" "3.75L + 1L + 1L = 5.75L"

The multiplier, "10^{-2}" or division by "100", is done to convert the given masses from grams to newtons.

moment about A "= 5.75L = F_{1} \times L"

"F_{1} = 5.75N"

sum of downward forces "= \frac{500N}{100} + \frac{200N}{100} + \frac{300N}{100}"

"5N + 2N + 3N = 10N"

Therefore, force, "F_{2}" on the other end B will be,

"F_{2} = 10N - 5.75N = 4.25N"